OpenStudy (anonymous):

Solve each equation by completing the square. Express all radicals to the nearest hundredth. x^2 - 10x + 16 = 0

OpenStudy (anonymous):

\[\eqalign{&x^2-10x+16=0 \\ &(x^2-10x)+16=0 \\ &(x^2-10x+25-25)+16=0 \\ &(x^2-10x+25)+16-25=0 \\ &(x-5)^2-9=0 \\ &(x-5)^2=9 \\ &\sqrt{(x-5)^2}=\pm\sqrt{9} \\ &x-5=\pm 3 \\ &x=\pm3+5 \\ &x=5\pm 3 \\ &(x=5+3) or(x=5-3) \\ &(x=8)or(x=2)\\}\]

OpenStudy (anonymous):

I have another problem... PLEASE RESPOND^^^

OpenStudy (anonymous):

x^2 - 12x +32 = 0